OpenStudy (anonymous):

A gym offers regular memberships for $80 per month and off-peak memberships for $60 per month. Last month, the gym sold a total of 420 memberships for a total of $31,100. How many memberships sold were regular memberships?

OpenStudy (bbypanda16):

your answer is 295

OpenStudy (bbypanda16):

regular memberships is x regular is y 80x+60y=31100 x+y=420 or you could just do 295*80=23600 420-295=125 125*60=7500 23600+7500=31100 so the answer is 295

OpenStudy (anonymous):

thankyou

OpenStudy (bbypanda16):

Anytime :)